# Semantic Communication and Control Co-Design for Multi-Objective Correlated Dynamics

### v1

@@ Abstract
This letter introduces a machine-learning approach to learning the semantic dynamics of correlated systems with different control rules and dynamics. By leveraging the Koopman operator in an autoencoder (AE) framework, the system’s state evolution is linearized in the latent space using a dynamic semantic Koopman (DSK) model, capturing the baseline semantic dynamics. Signal temporal logic (STL) is incorporated through a logical semantic Koopman (LSK) model to encode system-specific control rules. These models form the proposed logical Koopman AE framework that reduces communication costs while improving state prediction accuracy and control performance, showing a $91.65\%$ reduction in communication samples and significant performance gains in simulation.
Index Terms:
6G, signal temporal logic, predictive control, communication-control co-design.
@@ I Introduction
The rapid advancement of communication and control systems is foundational for advancing modern sixth-generation (6G) communication applications [1, 2]. However, these systems often operate in isolation, limiting scalability and complicating stability in distributed environments with constrained communication. To overcome these challenges, communication and control co-design (CoCoCo) approaches optimize communication resources without compromising control performance. For example, existing approaches focus on control scheduling and resource allocation, adapting to channel and control states for improved system stability [3, 4, 5]. While effective, these strategies prioritize data exchange over extracting meaningful dynamics, a crucial factor in ensuring control effectiveness.
In the meantime, recent machine learning (ML) advancements have introduced semantic and goal-oriented communication, particularly in beyond fifth generation (B5G) and 6G networks [6, 7, 8]. Semantic communication, aligned with CoCoCo’s objective, enhances communication efficiency and effectiveness by transmitting only task-relevant information. Unlike traditional methods, it leverages ML to extract and convey meaningful data rather than raw information. This approach suggests that so-called semantic CoCoCo will play a pivotal role in future systems, though challenges in communication cost, stability, and scalability across different systems persist [9]. While most semantic CoCoCo approaches rely on ML models that require significant training data, correlated control systems can exploit shared semantic structure, using a reference system’s dynamics. In contrast, systems with distinct behaviors often need independent models, posing scalability challenges.
Building on our preliminary work in [9], which demonstrated the effectiveness of controlling multi-objective identical systems, this paper addresses a new challenge: controlling multi-objective distinct control systems. These are distinct yet correlated mixed logical dynamical (MLD) systems, each with unique objectives and dynamics, but sharing baseline dynamics. These systems differ primarily in control rules and dynamic variations. We leverage a Koopman operator theory [10] within an AE architecture to linearize system state evolution in the latent space using a dynamic semantic Koopman (DSK) model, which captures the baseline semantic structure. To handle the system-specific variations, we incorporate signal temporal logic (STL) [11] into the latent space through the logical semantic Koopman (LSK) model to encode the system-specific control rules. Pre-training the encoder-decoder enables the AE to adapt to dynamic variations. Composing these models forms a novel semantic CoCoCo approach proposed in this letter as the logical Koopman AE framework.
To train the logical Koopman AE across multiple systems, we first train the encoder, decoder, and DSK model on a reference system devoid of control rules and dynamics variations, ensuring their applicability across different systems. Next, each system’s LSK model is trained by optimizing a loss function based on the local control rule. This approach significantly reduces communication costs associated with training each system’s dynamics, effectively addressing scalability without compromising control performance. Simulation on five distinct inverted cart-pole systems shows that our proposed framework yields a $75.78\%$ improvement in average state prediction capabilities while reducing communication samples by $91.65\%$ and improving the control performance by $92.91\%$ at a $15\,\mathrm{dB}$ signal-to-noise ratio (SNR) and two-dimensional state representations.
@@ II-A System Model
Consider controlling multi-objective distinct correlated control systems composed of plants, sensors, and remote controllers, as depicted in Fig. 1. The set of plants is denoted by $\mathcal{I}$, where each plant represents a non-linear dynamic process controlled by an actuator, with sensors sampling the process state. These sampled states are sent to a remote controller with high computational power, calculating the target control commands. The commands are then transmitted back to the actuators to drive the process to the desired behavior. The sensors and actuators are co-located and share a transceiver for wireless communication, while the remote controller communicates with the sensors via wireless channels. Each plant follows distinct control rules to meet objectives like stability and safety. Further details on the control and communication systems are provided next.
II-A1 Control System
Each plant’s sensor samples its $p$-dimensional state at a sampling rate $\tau_{o}$ and transmits these to the remote controller over a wireless channel. For plant $i\in\mathcal{I}$, the state at time $t=k\tau_{o}$ is given by $\mathbf{x}_{i,k}\in\mathbb{R}^{p}$. Upon receiving $\mathbf{x}_{i,k}$, the remote controller computes an optimal control command $\mathbf{u}_{i,k}\in\mathbb{R}^{q}$ and sends it to the actuator, which then influences the plant’s state. The process state evolves as a control-affine non-linear dynamic system as [12]
@@ II-A1 Control System
under given control rules $\varphi_{i}$, where $\mathbf{n}_{s,k}\in\mathbb{R}^{p}$ is system noise, modeled as independent and identically distributed (i.i.d.) Gaussian random vector. $\mathbf{f}^{s}_{\varphi_{i}}$ represents the non-linear state transition function and $\mathbf{F}^{u}_{\varphi_{i}}$ is the state-dependent control function, both are subject to predefined control rules $\varphi_{i}$. Each control rule, $\varphi_{i}$, will be formalized as an STL formula, translating natural language rules into a mathematical format. This concept and its integration into the logical Koopman AE architecture are further discussed in Section III.
Figure 1: An illustration of the multi-objective distinct control systems.
@@ II-A2 Wireless Communication
Each plant communicates with the remote controller using a dedicated orthogonal channel. Both uplink (sensor-to-controller) and downlink (controller-to-actuator) transmissions share the same channel, alternating via a time-division duplex. Channel reciprocity ensures consistent transmission power $P_{i}$ for both directions. Wireless channels are modeled using path loss and Rayleigh fading, with path loss defined as
@@
where $H_{i,k}$ is the Rayleigh fading channel gain, and $N_{c}$ is the power of additive white Gaussian noise. The channel capacity for plant $i$ at time $k$ is expressed as
@@ II-B Background Principles
The Koopman operator provides a linear representation of non-linear dynamics, enabling efficient state prediction for closed-loop control while reducing communication overhead. STL translates knowledge-based control rules into a formal mathematical framework. Brief overviews of both concepts are provided next.
@@ II-B1 Koopman Operator
Consider a real-valued measurement function $\psi:\mathbb{R}^{P}\rightarrow\mathbb{R}$, termed as an observable, residing within an infinite-dimensional Hilbert space. The Koopman operator $\mathcal{K}$, a linear operator on this space, evolves the observable such that [13]
@@
The Koopman operator linearizes non-linear dynamics in observable space but operates in an infinite-dimensional space, posing challenges. To address this, we identify an invariant subspace spanned by Koopman eigenfunctions ${\psi_{1},\psi_{2},\cdots,\psi_{d}}$, where $d$ is a finite positive integer, allowing for a finite-dimensional representation. Applying the Koopman operator to this subspace keeps the system in this subspace. Leveraging this linearity structure and linearly embedding control commands within observable space, we derive a global linear approximation of non-linear dynamics in (1) as [14]
@@
where $\Psi(\cdot)\in\mathbb{R}^{d}$ is a vector of Koopman eigenfunctions, $\mathbf{K}_{A}\in\mathbb{R}^{d\times d}$ is the state-based transition Koopman matrix, and $\mathbf{K}_{B}\in\mathbb{R}^{d\times q}$ is the control-based Koopman matrix. However, discovering Koopman eigenfunctions from finite data remains challenging, so we adopt an AE-based deep learning approach to address this, as detailed later.
@@ II-B2 Signal Temporal Logic
STL is a formal language for analyzing and interpreting discrete-time signals from control systems [15]. STL defines control rules over predicates, such as $\mu_{c}:=\mathbf{x}_{i}(p)<c$, where $c$ is a scalar. The syntax of an STL formula is given by
@@
where $a,b\in\mathbb{Z}_{+}$ is finite discrete-time bounds, with $0\leq a<b$, and $\varphi$ and $\vartheta$ are STL formulas. |||| (pipe symbol) separates between STL formulas, $\land$ (and) is boolean operator, and $\square$ (always) is temporal operator.
A timed trace signal $s_{i}$ records ordered sequence system state variables with timestamps. The satisfaction of an STL formula $\varphi$ by a timed trace signal $s_{i,k}$ is quantified by the robustness function $\rho$, which measures how strongly the signal satisfies or violates the formula: [15]
@@
The robustness function maps a timed trace and STL formula to a real value: positive for satisfaction and negative for violation. Using a parse tree for STL formulas, we generate a directed acyclic graph (DAG) to compute the robustness efficiently, with each sub-graph representing a node in the STL formula’s operation sequence [11, 9]. This resulting computation graph reflects the robustness of STL formulas as described in (II-B2).
III Logical Koopman Autoencoder
(a) DSK model.
(b) DSK $+$ LSK model.
Figure 2: The logical Koopman AE models for controlling two MLD systems.
This section introduces a novel semantic CoCoCo for controlling multi-objective distinct control systems, utilizing a logical Koopman AE inspired by the adapter technique in [16], which adapts a pre-trained model to new tasks without full retraining. In this framework, the Koopman model incorporates a trainable LSK model while keeping the DSK model frozen. The architecture consists of an encoder, a decoder, and a Koopman model with state and action components. The Koopman model shown in Fig. 2 is divided into four parts: two for the DSK model (state and action matrices) and two for the LSK model (state and action matrices), which govern the linear state evolution based on control rules. Training occurs in three phases: first, the encoder, decoder, and DSK model are trained on a reference system, then transferred to other systems where the LSK model is fine-tuned based on specific control rules. Finally, the systems are monitored remotely using linear predictions through the logical Koopman AE.
@@ III-A Dynamic Semantic Koopman Model
In the first phase of our remote control approach, a reference plant with simple dynamics and no specific control rule (denoted as $\varphi_{i}:=\mathsf{NULL}$) is selected as a benchmark. Here, we aim to optimize the encoder $\Psi(\cdot)$, decoder $\Psi^{-1}(\cdot)$, and DSK model, represented by matrices $\mathbf{D}_{A}$ and $\mathbf{D}_{B}$. The optimization problem is formulated as
${\footnotesize\underset{\mathbf{D}_{A},\mathbf{D}_{B},\Psi,\Psi^{-1}}{\arg\min}\;\frac{1}{K_{s}}\sum_{k=1}^{K_{s}}\left\lVert\mathbf{x}_{i,k}-\Psi^{-1}\left(\mathbf{D}_{A}\mathbf{z}_{i,k}+\mathbf{D}_{B}\mathbf{u}_{i,k}\right)\right\rVert^{2}_{2},}$ (10)
where $\mathbf{z}_{i,k}=\Psi(\mathbf{x}_{i,k})$ is the state representations of plant $i$ at time $k$. Due to the complexity of obtaining an analytical solution of (10) for long-term prediction, we leverage deep learning and define three loss functions defined as.
@@
    A reconstruction loss ($\mathcal{L}_{1}$) ensures accurate reconstruction of the system states and is defined as follows.
    $\mathcal{L}_{1}=\frac{1}{K_{s}}\sum_{k=1}^{K_{s}}||\mathbf{x}_{i,k}-\Psi^{-1}\left(\mathbf{z}_{i,k}\right)||^{2}_{2},$ (11)
    where $K_{s}$ is the number of system states observed in the first phase of remote controlling.
@@
    A linear dynamics loss ($\mathcal{L}_{2}$) enforces the linearity of system state evolution within the latent space over $k^{\prime}$ time steps, as shown as follows.
    $\mathcal{L}_{2}=\frac{1}{K_{s}}\sum_{k=1}^{K_{s}}||\mathbf{z}_{k+k^{\prime}}-\mathbf{D}_{A}^{k^{\prime}}\mathbf{z}_{i,k}+\mathbf{D}_{B}^{k^{\prime}}\mathbf{u}_{i,k}||^{2}_{2}.$ (12)
@@
    $\mathcal{L}_{3}=\frac{1}{K_{s}}\sum_{k=1}^{K_{s}}||\mathbf{x}_{k+k^{\prime}}-\Psi^{-1}(\mathbf{D}_{A}^{k^{\prime}}\mathbf{z}_{i,k}+\mathbf{D}_{B}^{k^{\prime}}\mathbf{u}_{i,k})||^{2}_{2}.$ (13)
The total loss function combines these three losses with an $\mathit{l}_{2}$ regularization term to prevent overfitting, leading to the final optimization problem formulated as.
@@
where $c_{n}$ are positive real coefficients and $\mathbf{W}$ represents the network weights. During this phase, frequent wireless transmission of system states and their representations is required to solve the optimization problem in (14).
In the second phase starting at time $t=k\tau_{o}$, after training the DSK model, the sensor transmits state representations $\mathbf{z}_{i,k}$ to the controller, which computes control commands $\mathbf{u}_{i,k}$. The controller predicts future states at time $t=(k+l)\tau_{o}$ using the DSK model and decoder as
@@
for $l\in\mathbb{Z}_{+}$. To counter prediction errors from communication noise and system uncertainties, periodic model fine-tuning is required, after which other plants adopt the trained models for their control phases outlined below.

### v2

# Semantic Communication and Control Co-Design for Multi-Objective Distinct Systems

@@ Abstract
This letter introduces a deep learning approach to predict the semantic control dynamics of different systems with different dynamics and control rules. By leveraging the Koopman operator in an autoencoder (AE) framework, the system’s state evolution is linearized in the latent space using a dynamic semantic Koopman (DSK) model, capturing the baseline semantic dynamics. The signal temporal logic (STL) is incorporated through a logical semantic Koopman (LSK) model to encode system-specific control rules. These models form the proposed semantic logical Koopman AE that reduces communication and computational costs without compromising control performance, showing a $91.65\%$ reduction in communication samples and significant performance gains in simulation.
@@ I Introduction
The rapid advancement of communication and control technologies underpins the development of emerging sixth generation (6G) applications, such as autonomous systems and industrial automation [1, 2]. However, traditional system architectures treat communication and control separately, leading to scalability bottlenecks and stability challenges in distributed and resource-constrained environments. To bridge this gap, communication and control co-design (CoCoCo) has emerged as a promising paradigm that optimizes communication resources without compromising control performance. Existing CoCoCo approaches primarily focus on adaptive scheduling and resource allocation based on channel and control states [3, 4, 5], yet they often overlook the underlying explicit system dynamics that govern multi-objective control systems and rely on lossy compression techniques to reduce communication overhead.
In parallel, the emergence of goal-oriented communication (GoC) and semantic communication (SC) has reshaped perspectives on information transmission for 6G networks. GoC prioritizes the transmission of task-relevant information required to achieve specific control or decision-making objectives, while SC focuses on extracting and transmitting meaningful representations from high-dimensional data that retain essential semantics of the information [6, 7, 8, 9]. Leveraging deep architectures such as auto-encoder (AE), SC enables the extraction and transmission of low-dimensional semantic representations, sufficient to accomplish tasks at the receiver. By transmitting only what is necessary to accomplish tasks, GoC directly enhances the effectiveness of achieving task-specific control objectives, while SC improves communication efficiency by eliminating redundant information. Building upon these, semantic CoCoCo emerges as an integrated concept that simultaneously benefits from both communication approaches: aligns with GoC to maximize task-specific control effectiveness and aligns with SC to minimize communication overhead. This synergy makes semantic CoCoCo particularly suitable for large-scale multi-objective distributed systems operating under limited communication and computational resources. Nonetheless, key challenges remain in jointly optimizing communication cost, control performance, and scalability across heterogeneous systems [10].
The multi-objective distinct control systems have been investigated in the literature. Existing approaches often employ control-aware scheduling and resource allocation to determine when and which control systems are granted transmission opportunities [11, 12, 13]. Specifically, these approaches prioritize control performance by dynamically allocating wireless resources based on control states or channel conditions, ensuring that the most critical control systems maintain acceptable performance under limited wireless resources. However, such approaches typically treat each control system separately rather than capturing cross-system interdependencies. More recently, multi-objective reinforcement learning has been explored to coordinate multiple distinct control tasks with shared wireless resources [14, 15]. It enables agents to learn policies that balance different objectives through reward shaping. Although it adapts to changing environments and jointly optimizes multiple objectives, it requires extensive interaction samples that lead to high communication and computational overhead, and it struggles to capture the underlying dynamics that couple among heterogeneous control systems.
Our prior work in [10] introduced a semantic CoCoCo framework for multi-objective control systems sharing the same dynamics but differing control rules, demonstrating that a correlated semantic structure significantly reduces communication overhead. To this end, we extend this approach to a more complex and realistic setting: multi-objective control of systems with both distinct dynamics and control rules. Specifically, we propose a novel semantic CoCoCo framework based on Koopman operator theory [16] for globally linearizing non-linear dynamics within latent space. Our proposed framework employs a dynamic semantics Koopman (DSK) model to extract a shared semantic structure from a reference system and a logic semantics Koopman (LSK) model to encode system-specific control rules using signal temporal logic (STL) [17]. The AE is pre-trained on the reference system and then fine-tuned using a small subset of data from each system to adapt to local dynamic variations. The composition of DSK and LSK forms the proposed semantic logical Koopman AE, a scalable, task-driven approach that preserves control performance while reducing communication costs.
Unlike [10], which is based on systems with identical system dynamics, the proposed approach addresses the scalability challenges in heterogeneous systems, enabling knowledge transfer across structurally distinct systems governed by logical control rules. Each system’s LSK model is trained using a temporal logic-based loss, allowing control rules to be embedded directly into latent dynamics. This design eliminates the need to train a full Koopman model per system, reducing computational and communication overhead. Extensive simulations on five distinct cart-pole systems validate the proposed approach, showing a $75.78\%$ improvement in average state prediction, a $91.65\%$ reduction in communication samples, and a $92.91\%$ increase in control performance under $15\,\mathrm{dB}$ signal-to-noise ratio (SNR) and using two-dimensional state representations.
@@ II-A System Model
Consider controlling multi-objective distinct control systems composed of plants, sensors, and remote controllers, as depicted in Fig. 1. The set of plants is denoted by $\mathcal{I}$, where each plant represents a non-linear dynamic process controlled by an actuator, with sensors sampling the process state. These sampled states are sent to a remote controller with high computational power to calculate the target control commands. The commands are then transmitted back to the actuators to drive the process to the desired behavior. The sensors and actuators are co-located and share a transceiver for wireless communication, while the remote controllers communicate with the sensors via wireless channels. Each plant follows different control rules to meet objectives such as stability and safety. Further details on the control and communication systems are provided next.
II-A1 Control Model
Each plant sensor samples its $p$-dimensional state at a sampling rate $\tau_{o}$ and transmits these to the remote controller over a wireless channel. For plant $i\in\mathcal{I}$, the state at time $t=k\tau_{o}$ is given by $\mathbf{x}_{i,k}\in\mathbb{R}^{p}$. Upon receiving $\mathbf{x}_{i,k}$, the remote controller computes an optimal control command $\mathbf{u}_{i,k}\in\mathbb{R}^{q}$ and sends it to the actuator, which then influences the plant state. The process state evolves as a control-affine non-linear dynamics as [18]
@@ II-A1 Control Model
under given control rules $\varphi_{i}$, where $\mathbf{n}_{s,k}\in\mathbb{R}^{p}$ is system noise, modeled as independent and identically distributed (i.i.d.) Gaussian random vector. $\mathbf{f}^{s}_{\varphi_{i}}$ represents the non-linear state transition function and $\mathbf{F}^{u}_{\varphi_{i}}$ is the state-dependent control function, both are subject to predefined control rules $\varphi_{i}$. Each control rule, $\varphi_{i}$, will be formalized as an STL formula, translating natural language rules into a mathematical format. This concept and its integration into the semantic logical Koopman AE are discussed further in Section III.
Figure 1: Illustration of multi-objective distinct and distributed control systems.
@@ II-A2 Wireless Communication
Each plant communicates with the remote controller using a dedicated orthogonal channel. Both uplink (sensor-to-controller) and downlink (controller-to-actuator) transmissions share the same channel, alternating via a time-division duplex. The channel reciprocity ensures consistent transmission power $P_{i}$ in both directions. Wireless channels are modeled using path loss and Rayleigh fading, with path loss defined as
@@
where $H_{i,k}$ is the Rayleigh fading channel, and $N_{c}$ is the power of additive white Gaussian noise. The channel capacity for plant $i$ at time $k$ is expressed as
@@ II-B Background Principles
The Koopman operator provides a global linear representation of non-linear dynamics, enabling efficient state prediction for closed-loop control. The STL translates natural language-based control rules into a formal mathematical form. Brief overviews of both concepts are provided below.
@@ II-B1 Koopman Operator
Consider a real-valued measurement function $\psi:\mathbb{R}^{P}\rightarrow\mathbb{R}$, termed as an observable, residing within an infinite-dimensional Hilbert space. The Koopman operator $\mathcal{K}$, a linear operator in this space, evolves the observable such that [19]
@@
The Koopman operator linearizes non-linear dynamics in an observable space but operates in an infinite-dimensional space, posing challenges. To address this, we identify an invariant subspace spanned by Koopman eigenfunctions $\psi_{1},\psi_{2},\cdots,\psi_{d}$}, where $d$ is a finite positive integer, allowing for a finite-dimensional representation. Applying the Koopman operator to this subspace keeps the system in this subspace. Leveraging this linearity structure and linearly embedding control commands within observable space, we derive a global linear approximation of non-linear dynamics in (1) as [20]
@@
where $\Psi(\cdot)\in\mathbb{R}^{d}$ is a vector of Koopman eigenfunctions, $\mathbf{K}_{A}\in\mathbb{R}^{d\times d}$ is the state-based transition Koopman matrix, and $\mathbf{K}_{B}\in\mathbb{R}^{d\times q}$ is the control-based Koopman matrix. Although the Koopman operator allows global linear representations of non-linear dynamics within the latent space to utilize linear control and prediction techniques, discovering Koopman eigenfunctions from finite data remains challenging. To address this, we adopt a AE-based deep learning approach among various deep learning architectures, such as AE and transformers, due to its simplicity and computational efficiency in learning lower-dimensional representations of system states.
@@ II-B2 Signal Temporal Logic
STL is a formal mathematical language to specify the temporal properties of discrete-time signals in control systems [21]. It enables the expression of time-bounded safety and control properties, making it suitable to guide the behavior of control systems under dynamic, real-time constraints. The STL defines control rules over predicates such as $\mu_{c}:=\mathbf{x}_{i}(p)<c$, where $c$ is a scalar. The syntax of a STL formula is given by
@@
where $a,b\in\mathbb{Z}_{+}$ are finite discrete-time bounds, with $0\leq a<b$, and $\varphi$ and $\vartheta$ are STL formulas. $|$ (pipe symbol) separates between STL formulas, $\land$ (and) is the boolean operator, and $\square$ (always) is the temporal operator that specifies the temporal properties of signals within the time interval $\left[a,b\right]$.
A timed trace signal $s_{i,k}$ records ordered sequence system state variables with timestamps. The satisfaction of an STL formula $\varphi$ by a timed trace signal $s_{i,k}$ is quantified by the robustness function $\rho$, which measures how strongly the signal satisfies or violates the formula: [21]
@@
The robustness function maps a timed trace and STL formula to a real value: positive for satisfaction and negative for violation. Using a parse tree for STL formulas, we generate a directed acyclic graph to compute the robustness efficiently, with each sub-graph representing a node in the STL formula’s operation sequence [17, 10]. This resulting computation graph reflects the robustness of STL formulas as described in (II-B2).
III Semantic Logical Koopman Auto-encoder
(a) First phase for training the encoder, decoder, and DSK model.
(b) Second phase for training the LSK model.
Figure 2: Training procedure of the semantic logical Koopman AE for controlling two MLD systems with two main phases.
This section introduces a novel semantic CoCoCo for controlling multi-objective distinct control systems, utilizing a semantic logical Koopman AE inspired by the adapter technique in [22], which adapts a pre-trained model to new tasks without full retraining. In this framework, the Koopman model incorporates a trainable LSK model while keeping the DSK model frozen. The architecture consists of an encoder, a decoder, and a Koopman model with state and action components. The Koopman model shown in Fig. 2 is divided into four parts: two for the DSK model (state and action matrices) and two for the LSK model (state and action matrices), which govern the linear state evolution based on control rules. The training procedure consists of two phases. In the first phase, the encoder, decoder, and DSK model are jointly trained on a reference system, then, in the second phase, these components are transferred to other systems where the AE is fine-tuned using a small subset of local data to adapt to system-specific dynamics. Concurrently, the LSK model is trained to encode the local control rules via a temporal logic-based loss. Once trained, each system is monitored remotely through efficient linear predictions by the semantic logical Koopman AE.
@@ III-A Dynamic Semantic Koopman Model
In the first phase of our remote control approach, a reference plant with simple dynamics and no specific control rule (denoted as $\varphi_{i}:=\mathsf{NULL}$) is selected as a benchmark. Here, we aim to optimize the encoder $\Psi(\cdot)$, decoder $\Psi^{-1}(\cdot)$, and DSK model, represented by the matrices $\mathbf{D}_{A}$ and $\mathbf{D}_{B}$. The optimization problem is formulated as.
${\footnotesize\underset{\mathbf{D}_{A},\mathbf{D}_{B},\Psi,\Psi^{-1}}{\arg\min}\;\frac{1}{K_{s}}\sum_{k=1}^{K_{s}}\left\lVert\mathbf{x}_{i,k+k^{\prime}}-\Psi^{-1}\left(\tilde{\mathbf{z}}_{i,k+k^{\prime}}\right)\right\rVert^{2}_{2},}$ (10)
with the future predicted state representation of the reference plant $i$ at time $k+k^{\prime}$ given as.
$\tilde{\mathbf{z}}_{i,k+k^{\prime}}=\mathbf{D}^{k^{\prime}}_{A}\mathbf{z}_{i,k}+\mathbf{D}^{k^{\prime}}_{B}\mathbf{u}_{i,k},$ (11)
where $\mathbf{z}_{i,k}=\Psi(\mathbf{x}_{i,k})$ is the state representations of plant $i$ at time $k$. Due to the complexity of obtaining an analytical solution of (10) for long-term prediction, we leverage deep learning and define three loss functions given as.
@@
    A reconstruction loss ($\mathcal{L}_{1}$) ensures accurate reconstruction of the system states and is defined as.
[…]
    where $K_{s}$ is the number of system states observed in the first phase of remote control.
@@
    A linear dynamics loss ($\mathcal{L}_{2}$) enforces the linearity of system state evolution within the latent space, as follows.
[…]
    where $\tilde{\mathbf{z}}_{i,k+k^{\prime}}=\Psi\left(\mathbf{x}_{i,k+k^{\prime}}\right)$ is the future state representation of plant $i$ at time $k+k^{\prime}$.
@@
[…]
The total loss function combines these three losses with a $\mathit{l}_{2}$ regularization term to prevent overfitting, leading to the final optimization problem formulated as.
@@
where $c_{n}$ are positive real coefficients and $\mathbf{W}$ represents the model weights. During this phase, frequent wireless transmission of system states and their representations is required to solve the optimization problem in (15).
In the second phase, starting at time $t=k\tau_{o}$, after training the DSK model, the sensor transmits state representations $\mathbf{z}_{i,k}$ to the remote controller, which computes control commands $\mathbf{u}_{i,k}$. The controller predicts future states at time $t=(k+l)\tau_{o}$ using the DSK model and decoder as
@@
for $l\in\mathbb{Z}_{+}$. To counter prediction errors from communication noise and system uncertainties, frequent model fine-tuning is required, after which other plants adopt the trained models for training their own models, outlined below.